OpenStudy (anonymous):

(-4x + 2) (-2) + y + (-x)

OpenStudy (anonymous):

(-4x + 2) (-2) + y + (-x) = (8x - 4) + y + (-x) = 8x - 4 + y - x = 8x - x -4 + y =7x - 4 + y =7x + y - 4